# An average Brun–Titchmarsh theorem and shifted primes with a large prime factor

### v2

@@ 1. Introduction
In 1996, Baker and Harman [1] presented a mixed approach to this problem. They combined the traditional average Brun–Titchmarsh approach (with a refined version of [[12], Théorème 3] which will be used in Section 6.10) with the alternative sieve developed by Harman [15], using the estimates of Bombieri–Friedlander–Iwaniec type (see [5] and [6], for example) as arithmetical information inputs. Their method is more powerful when $\theta$ is near to $\frac{1}{2}$. In this way they got $\boldsymbol{\Theta}=0.676$. Two years later, they [2] further developed a variant of the three-dimensional alternative sieve used in [1] and obtained a saving over one-dimensional sums, leading to a better result $\boldsymbol{\Theta}=0.677$. Although this result only improved the previous one by $0.001$, it was still a improvement since they used a continuous version of alternative sieve, and any saving on a single point became quite small after integrating. Because of the lack of new arithmetical information, even $0.678$ seems impossible without any new things.
Recently, Maynard [28] got a new breakthrough on the distribution of primes in arithmetical progressions to large moduli $q=q_{1}q_{2}\sim x^{\frac{11}{21}-\varepsilon}$. The main theorem in his paper is not applicable here, but we find that one of his lemmas (Proposition 8.3) can be used here as a new arithmetical information input. By inserting this into the sieving process in [1] and [2], together with a careful numerical calculation for sieve integrals, we get the following sharper result.
@@
Throughout this paper, we always suppose that $\varepsilon$ is a sufficiently small positive constant and $x$ is sufficiently large. Let $\theta\in[0.5,0.679]$ and $Q=x^{\theta}$. The letter $p$, with or without subscript, is reserved for prime numbers. We define the boolean function as
$\texttt{Boole}(\textbf{X})=\begin{cases}1&\text{if }\textbf{X}\text{ is true},\\ 0&\text{if }\textbf{X}\text{ is false}.\end{cases}$ (5)
@@ 2. An outline of the proof
In the process of giving asymptotic formulas like (13) in Harman’s sieve, one needs the following types of arithmetical information (with all variables are dyadic constrained, and all coefficients are divisor-bounded):
@@
3. Type-II arithmetical information
@@
$\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}\cdots m_{j}\in\mathcal{A}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant j\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}}=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}\cdots m_{j}\in\mathcal{B}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant j\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}}.$
The three conditions we want the coefficients to satisfy are as follows:
@@ 6. The final decomposition
$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{101}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{104}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{105}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{106}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{107}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{108}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
@@
$\left(\int_{\kappa}^{\frac{\theta}{2}}\int_{1-\theta-t_{1}}^{\theta-t_{1}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{201}]\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{204}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{205}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in L_{206}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{t_{3}}^{\frac{1-t_{1}-t_{2}-t_{3}}{2}}\int_{t_{4}}^{\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{207}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\int_{\kappa}^{\min\left(t_{3},\frac{1-t_{1}-t_{2}-t_{3}}{2}\right)}\int_{\kappa}^{\min\left(t_{4},\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{2}\right)}\right$.
$\left.\quad\quad\quad\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{208}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{4-7\theta}^{\frac{3}{7}}\texttt{Boole}[(t_{1})\in L_{209}]\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\texttt{Boole}[(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\min\left(t_{2},\frac{1-t_{1}-t_{2}}{2}\right)}\texttt{Boole}[(t_{1},t_{2},t_{3})\in L_{212}]\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$

### v3

@@ 1. Introduction
In 1996, Baker and Harman [1] presented a mixed approach to this problem. They combined the traditional average Brun–Titchmarsh approach (with a refined version of [[12], Théorème 3] which will be used in Section 6.10) with the alternative sieve developed by Harman [15], using the estimates of Bombieri–Friedlander–Iwaniec type (see [5] and [6], for example) as arithmetic information inputs. Their method is more powerful when $\theta$ is near to $\frac{1}{2}$. In this way they got $\boldsymbol{\Theta}=0.676$. Two years later, they [2] further developed a variant of the three-dimensional alternative sieve used in [1] and obtained a saving over one-dimensional sums, leading to a better result $\boldsymbol{\Theta}=0.677$. Although this result only improved the previous one by $0.001$, it was still a improvement since they used a continuous version of alternative sieve, and any saving on a single point became quite small after integrating. Because of the lack of new arithmetic information, even $0.678$ seems impossible without any new things.
Recently, Maynard [28] got a new breakthrough on the distribution of primes in arithmetic progressions to large moduli $q=q_{1}q_{2}\sim x^{\frac{11}{21}-\varepsilon}$. The main theorem in his paper is not applicable here, but we find that one of his lemmas (Proposition 8.3) can be used here as a new arithmetic information input. By inserting this into the sieving process in [1] and [2], together with a careful numerical calculation for sieve integrals, we get the following sharper result.
@@
Throughout this paper, we always suppose that $\varepsilon$ is a sufficiently small positive constant and $x$ is sufficiently large. Let $\theta\in[0.5,0.679]$ and $Q=x^{\theta}$. The letter $p$, with or without subscript, is reserved for prime numbers.
@@ 2. An outline of the proof
In the process of giving asymptotic formulas like (13) in Harman’s sieve, one needs the following types of arithmetic information (with all variables are dyadic constrained, and all coefficients are divisor-bounded):
@@
3. Type-II arithmetic information
@@
$\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}\cdots m_{j}\in\mathcal{A}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant j\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}}=(1+o(1))\frac{1}{\varphi(q)}\sum_{q\sim Q}\gamma_{q}\sum_{\begin{subarray}{c}m_{1}\cdots m_{j}\in\mathcal{B}^{q}\\ m_{i}\sim M_{i}\\ 1\leqslant i\leqslant j\end{subarray}}a_{1,m_{1}}a_{2,m_{2}}\cdots a_{j,m_{j}}$
(The errors, here and in lemmas below, can be bounded by $x(\log x)^{-A}$ for some large $A>1$.) The three conditions we want the coefficients to satisfy are as follows:
@@ 6. The final decomposition
$\left(\int_{(t_{1},t_{2},t_{3})\in L_{101}}\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{104}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{105}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{106}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{107}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{108}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
@@
$\left(\int_{(t_{1},t_{2},t_{3})\in L_{201}}\frac{\omega\left(\frac{t_{2}}{t_{1}}\right)\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}^{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{1}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{2}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{204}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{205}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4})\in L_{206}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{t_{4}}\right)}{t_{1}t_{2}t_{3}t_{4}^{2}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
- $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{207}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3},t_{4},t_{5})\in L_{208}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}-t_{5}}{t_{5}}\right)}{t_{1}t_{2}t_{3}t_{4}t_{5}^{2}}dt_{5}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1})\in L_{209}}\frac{\omega\left(\frac{1-t_{1}}{t_{1}}\right)}{t_{1}^{2}}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3})\in\boldsymbol{D}_{3}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{\kappa}\right)}{t_{1}t_{2}t_{3}}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\frac{1}{\kappa}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\int_{\kappa}^{\frac{1}{2}}\mathbbm{1}_{(t_{1},t_{2},t_{3},t_{4})\in\boldsymbol{D}_{4}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}-t_{4}}{\kappa}\right)}{t_{1}t_{2}t_{3}t_{4}}dt_{4}dt_{3}dt_{2}dt_{1}\right)$
$+$ $\left(\int_{(t_{1},t_{2},t_{3})\in L_{212}}\frac{\omega\left(\frac{1-t_{1}-t_{2}-t_{3}}{t_{3}}\right)}{t_{1}t_{2}t_{3}^{2}}dt_{3}dt_{2}dt_{1}\right)$